Bill is 7 years older than his dog. The sum of their ages is 19. How old is Bill?

Answer by nyc_function(2741)

If Bill's age is x, his dog's age is (x - 7)
That's how I would have done it, anyway. That way you can solve for Bill's age as x, which is what the question asks.
Now since the sum of their ages is 19, you know that:
x + (x - 7) = 19
2x - 7 = 19
2x = 19 + 7
2x = 26
x = 26/2
x = 13
You actually COULD get the answer with your equation, but x would be the dog's age.
2x + 7 = 19
2x = 19 - 7
2x = 12
x = 6
Notice that 6 is seven years younger than 13.

ANSWER

Bill is 13.